OpenStudy (anonymous):

Mike repairs televisions. His revenue equals R(h)=25+30h dollars for every h hours he spends repairing television. His overhead cost equals C(h)=5h2-10 dollars. After how many hours does he break even?

OpenStudy (anonymous):

Are you allowed to graph this, or do you need to solve it out algebraically?

OpenStudy (anonymous):

Also, is the 2 in 5h2 supposed to be an exponent?

OpenStudy (anonymous):

well, i don't know

OpenStudy (anonymous):

We'll do it algebraically then since that's what most teachers prefer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

And the question on the 5h2?

OpenStudy (anonymous):

Is the 2 a exponent?

OpenStudy (anonymous):

oh yes it is

OpenStudy (anonymous):

Okay then, here's what we need to do...The point where it breaks even is going to be found with profit=0 or when the revenue is the same as the cost. Do you know how to write a function for that?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

whoops sorry

OpenStudy (anonymous):

It's alright. Mine keeps going in and out as well

OpenStudy (anonymous):

Let me ask you this...Breaking even would be at a point where the worker receives no profit nor does he lose any money. Do you agree with that?

OpenStudy (anonymous):

Why is that?

OpenStudy (anonymous):

Because, he can't break even, if he did or didnt lose money wouldnt possibility be correct

OpenStudy (anonymous):

Well lets consider this...The definition of breaking even is "to reach a point in a business venture when the profits are equal to the costs". Do you agree that 1 + 1 = 0?

OpenStudy (anonymous):

no 1+1 does not equal 0

OpenStudy (anonymous):

Lol So sorry. I meant 1-1=0

OpenStudy (anonymous):

We both know that 1-1=0. With that being said, if our "break even point" is where the cost equals the profit, then we would receive no profit, nor would we lose any money. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So back to what I said earlier, "Breaking even would be at a point where the worker receives no profit nor does he lose any money." Do you understand why that is actually correct?

OpenStudy (anonymous):

If you don't, just let me know.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Okay as in you don't?

OpenStudy (anonymous):

no i do

OpenStudy (anonymous):

I'm trying to figure out how many hours is asking me

OpenStudy (anonymous):

Okay. Excellent. Now that we know that our cost and revenue must be the same in order to reach our break even point, we know that (Revenue) - (Cost) = 0 So R(h) - C(h) = 0 (25+30h) - (5h^2-10) = 0

OpenStudy (anonymous):

okay,so we have that just how are we going to solve it?

OpenStudy (anonymous):

(25+30h) - (5h^2-10) = 0 Simplify and get... -5h^2 + 30h + 35 = 0 Make sense so far?

OpenStudy (anonymous):

Don't worry, I'm getting there.

OpenStudy (anonymous):

yes, it makes sense

OpenStudy (anonymous):

Okay, no we are going to use the quadratic formula to solve for the 0's in our function... \[h=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[h= \frac{ −(30)±\sqrt{(30)^2−4(-5)(35)}}{ 2(-5)}\]

OpenStudy (anonymous):

Can you solve that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Go for it. :)

OpenStudy (anonymous):

-10 +/- ^200/-10

OpenStudy (anonymous):

Um, that's not what I got. Let me check again.

OpenStudy (anonymous):

yea, i think idk

OpenStudy (anonymous):

\[h=\frac{ −(30)±\sqrt{(30)^2−4(−5)(35)} }{ 2(−5) }\] \[h=\frac{ −30±\sqrt{30^2-700} }{ -10 }\] \[h=\frac{ −30±\sqrt{900-700} }{ -10 }\] \[h=\frac{ −30±\sqrt{1600} }{ -10 }\] \[h=\frac{ −30±40 }{ -10 }\]

OpenStudy (anonymous):

Just a moment...

OpenStudy (anonymous):

My power is giving out here. There's a bad storm happening

OpenStudy (anonymous):

uh oh, not here until tomorrow, cause i'm in Florida

OpenStudy (anonymous):

\[h=\frac{−30±40}{−10}\] Because there's a ± sign, we can solve this 2 ways... \[h=\frac{−30-(40)}{−10}\] which -30 - 40 = -70 and -70/-10 = 7 so one of your zeroes is h=7 The other is... \[h=\frac{−30+(40)}{−10}\]which -30 + 40 = 10 and 10/-10 = -1 so your other zero is h=-1

OpenStudy (anonymous):

Can you tell which zero (aka the amount of hours) is the correct answer for this problem?

OpenStudy (anonymous):

i would with 7h

OpenStudy (anonymous):

Correct. You can't have a negative amount of hours, so -1 makes no sense. Therefore, 7 hours would be your break even point.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

You can see this on a graph if you go to www.desmos.com/calculator, and in one of the spaces put (25+30x)-(5x^2-10). You'll see that -1 and 7 are the x-intercepts.

OpenStudy (anonymous):

You're very welcome! :D

OpenStudy (anonymous):

well i gtg bye bye and thanks for helping me

OpenStudy (anonymous):

You're welcome and goodnight.